How do i find the molar mass of this metal? the mass is .079 g, the volume of gas in the buret is 30.2 ml, the water height differential is 429 mm, the temperature of the gas is 22 C, and the barometric pressure is 761 mm Hg.
Answer #1

Given the mass of metal taken, W = 0.079 g

temperature of the gas, T = 22 DegC = 22 + 273 = 295 K

Volume of gas, V = 30.2 mL = 30.2 mL x (1L / 1000 mL) = 0.0302 L

Given water height differential, P(liquid water) = 429 mm x (0.99751 / 13.596) = 31.475 mm Hg

Aqueous vapor pressure due to water vapor, P(water vapor) = 21.0 mm Hg

Hence pressure due to H2 gas, P = 761 mmHg - P(liquid water) - P(water vapor) =

=> P = 761 mmHg - 31.475 mm Hg - 21.0 mmHg = 708.525 mmHg = 0.9323 atm

Now we can calculate molar mass by applying ideal gas equation as

PV = nRT = WRT / M [Since n = W/M]

=> M = WRT / PV = (0.079 g x 0.0821 L.atm.mol-1K-1 x 295 K) / (0.9323 atm x 0.0302 L)

=> M = 67.96 g/mol (answer)